# Why do explanations fail? A typology and discussion on failures in XAI

### v1

@@ Introduction
Mismatch
First, we propose to define mismatched explanations when the format of the extracted information does not meet the users’ expectations towards the ML system (i.e., ”This is not what I want” in Fig. 2).
@@
Depending on the context of the interaction and the nature of the decision model’s outputs, users have been shown to have different needs and questions regarding the ML system and the user questions may also vary depending on the model’s output  (Liao, Gruen, and Miller 2020; van der Waa et al. 2021). In parallel, there is a huge diversity in the forms explanations can take (see e.g. (Dwivedi et al. 2023)), but explanation techniques do not necessarily take into account the context in which the explanations is sought by a user (Matarese, Rea, and Sciutti 2021). As a consequence, mismatches between the user questions and the generated explanations can occur, on the explanation type (e.g., feature importance or rules), locality (e.g., local or global), goal (e.g., factual or causal) and on the information complexity (e.g., expressed for ML practitioners or lay users), to name a few.
Such mismatches reflect the lack of user-centricity in the conception of the XAI system. Actually, most XAI approaches are designed without evaluating whether the explanations satisfy the needs of real users (Miller 2019). Instead, other criteria are used to evaluate the relevance of an XAI approach, such as the visual aspect of the explanations, its popularity or the ease of implementation (Mohseni, Zarei, and Ragan 2018; Barredo Arrieta et al. 2020). Such evaluations can bring to light that they fail to meet these needs: for example, during the co-design workshop for an AI-based diagnosis tool (Wang et al. 2019), many interviewed doctors reported that they would prefer alternative hypotheses (e.g., counterfactual examples) rather than factual explanations (e.g., local feature importance technique) that were initially implemented in the offered system.
@@
Mismatched explanations can lead to dissatisfaction and a reject of the overall ML system: It has been demonstrated that differences between initial expectations and actual experiences can affect both the user satisfaction and acceptance of a system (Bhattacherjee 2001). Along the same lines, the complexity of some explanation types can hinder users to understand the explanations (e.g., if they lack skills in AI), and thus the adoption of the whole ML system (see e.g. (Kaur et al. 2020) in the case of SHAP and GAM). It has also been demonstrated that users assign artificial agents human-like traits, and hence expect these agents to provide explanations using the same conceptual framework they are used to (De Graaf and Malle 2017). Yet, most XAI solutions proposed so far have not reached such a complete human-centred approach (Miller 2019; Matarese, Rea, and Sciutti 2021).
@@
Mitigating the mismatched explanations somehow obviously relies on applying user-centered approaches when designing XAI systems: two examples include adjusting the types of explanations according to users’ expectations to improve their satisfaction and acceptance (Riveiro and Thill 2022) or integrating a lexical alignment step to improve the understanding of explanations provided by a conversational agent (Srivastava, Theune, and Catala 2023). Another type of solution aims at integrating social and cognitive science’s theories in a theoretical framework for XAI system, to provide personalized and contextualized explanations (Matarese, Rea, and Sciutti 2021).
Counterintuitive explanations
We then propose to identify counterintuitive explanations when the explanation provided by the ML system is in contradiction with the prior knowledge of expert users (i.e., AI practitioners and domain experts). As compared to the mismatch failure for which explanations can be contradictory with users expectations in terms of format, counterintuitive failures can be observed when the content of the explanation itself is contradictory with users’ knowledge.
@@
Prior knowledge may take various definitions (Dochy and Alexander 1995), we view it here as ”stored knowledge about the world that have been acquired by an individual” (Brod, Werkle-Bergner, and Shing 2013), including domain expertise and past experience. Now in an XAI context, some users can find faithful explanations contradicting with such prior knowledge, leading them to perceive these explanations as different from what they anticipated (”This is not what I expected” in Fig. 2). This is all the more likely to happen as explanations are especially requested when the model’s output is perceived by users as abnormal or absurd (Riveiro and Thill 2022). In the same vein, past experience has been shown to lead to disagreement with the explanations (Sohn et al. 2019; Suffian et al. 2022).
@@
Counterintuitive explanations do not necessarily represent a failure: explanations that do not match user expectations can indeed be used to fix potential issues within the ML system (Kaur et al. 2020) or for knowledge discovery (Jiménez-Luna, Grisoni, and Schneider 2020). In other situations however, perceiving the provided explanations as counterintuitive can lead users to question the reliability of the prediction even when it is accurate (e.g., (Collaris, Vink, and van Wijk 2018) in the context of insurance fraud and (Palaniyappan Velumani et al. 2022)in the context of environmental studies), or of the ML system as a whole (e.g. (Nourani et al. 2021)). They can even perceive the explanations as being absurd if the latter do not correspond to their past experiences (Sohn et al. 2019). This aligns with findings from works in social sciences, which have shown that people tend to ignore an information that is inconsistent with their beliefs from past experiences (Thagard 1989; Nickerson 1998).
@@
Mitigating these failures requires to better align ML explanations and users’ prior knowledge. Some works in XAI propose to integrate the users’ knowledge directly in the explanation generation (e.g., (Jeyasothy et al. 2022)). Conversely, a different approach proposes to change the predictive model itself to facilitate the generation explanations that are aligned with users’ prior knowledge (Rieger et al. 2020). In a different perspective, another work proposes to co-design explanation interfaces together with expert in co-design workshops so as to integrate both their needs and knowledge (Wang et al. 2019).
Biased inferences
Finally, we propose to identify explanation failures relying on biased inferences when users make inaccurate interpretations of the explanations, due to cognitive biases (see Fig. 2). As compared to counterintuitive failures that occur for expert users, for biased inferences we consider mainly lay users that do not have expert knowledge nor past experiences. Yet, such failures can occur for any type of users as cognitive biases are inherent to all humans.
@@
Similarly to prior beliefs, biases can influence how users respond to different styles of explanations (Liao, Gruen, and Miller 2020; Miller 2019), leading eventually to their inaccurate interpretation (Bertrand et al. 2022; Nourani et al. 2021; Wang et al. 2019). We discuss below some common cognitive biases with their consequences on the interpretation of explanations.
@@
Several cognitive biases have been shown to trigger inaccurate interpretations of explanations (Bertrand et al. 2022). First, some biases can trigger over reliance in ML explanations. It has been demonstrated that having an explanation, regardless of its quality, increases trust (Eiband et al. 2019; Lai and Tan 2019). In other examples, it is shown that longer explanations are found to be more plausible than shorter ones (Fürnkranz, Kliegr, and Paulheim 2020), and that users may believe they understand better than what they actually do (Rozenblit and Keil 2002; Mueller et al. 2019; Chromik et al. 2021). On the other hand, other biases can trigger under reliance. For instance, the ”negativity bias” can cause lay users, in particular, to pay more attention or overweight negative information over positive one of the same strength (Kliegr, Bahník, and Fürnkranz 2021). It may lead users to pay more attention to negative outcomes of the ML system, thus eroding their trust (Pratto and John 1991). It has been demonstrated that showing the weaknesses of the system (e.g., competing explanations) or negative outcomes (e.g., such as a malignant diagnosis) early on can have a major influence on trust (Nourani et al. 2021). Finally, some other biases can trigger users to misapply the explanations: e.g., the ”insensitivity to sample size” bias may lead lay users to ignore the statistical significance of a statement (Fürnkranz, Kliegr, and Paulheim 2020); e.g., the ”availability” may lead lay users to believe that examples and events that easily come to mind are more representative than is actually the case (Wang et al. 2019); e.g., the ”primacy effect” bias may lead them to form an opinion based solely on the first piece of information received (Nourani et al. 2021).
@@
Discussion and research agenda
The typology of explanation failures presented in the previous section allows to understand why failures happen, how to mitigate them, and how to distinguish them from one another. In this section, we leverage this typology to discuss some key issues, and identify some promising research directions.
Towards a holistic XAI approach
@@
Many of the explanation failures discussed in the previous sections can be diagnosed as stemming from one of these components (model, explainer and user): they can for instance be due to a technical problem with the system. However, our analysis in the previous section also underlines that some errors actually arise from the interplay between the different components considered, rendering them incompatible: we studied the interaction between the explainer and the user in the third section about user-specific failures and discussed in the fourth section some cases of problematic interaction between the model and the explainer. For instance, issues like unstability may highlight a mismatch between the decision model’s behavior (volatile, local), the implicit assumptions made during explanation design (stable boundaries) and the similarity between instances perceived by the user.
@@
In the case of these interaction between components, the failure is not caused by a deficiency in either component, but rather from their misalignment. Consequently, some technical solutions to address these failures suggest adapting or replacing one of the components afterwards, sometimes taking a paradoxical turn: instead of questioning the explainer when observing a failure, it is sometime envisaged instead to train a new AI decision model that would be more adapted to this explainer. As an illustrative example, it has been proposed to build models with more stable behavior to mitigate instability issues of gradient-based explainer (Alvarez-Melis and Jaakkola 2018b), or models constrained to minimize the disagreement between LIME and SHAP (Schwarzschild et al. 2023). This may seem surprising, as explainers are usually leveraged to generate insights about the model, not the other way around.
@@
This underscores a significant challenge: the interplay between the elements of the system should be taken into account from the system’s inception. The three components of the system should thus be regarded as interconnected, rather than being designed independently. This does not mean abandoning post-hoc methods but rather anticipating their integration in the overall explanation process.
Multiple calls for a user-centric approach of XAI have been made, pushing for taking the user needs into account from the inception of the system and guiding the design of the XAI method (Wang et al. 2019). Nevertheless, more efforts should be pursued in studies on the interaction between AI models and explainers. One possibility is to draw inspiration from research in Integrative Design for software systems (Tumer and Smidts 2010) which proposes holistic design strategies for software systems. Design and monitoring of AI systems should thus be conceived in a holistic way, with any choice or change in the decision model prompting a reassessment of its compatibility with the explainer, and vice versa; and similarly for changes in user requirements.
Towards less one-size-fits-all and more transparent XAI solutions
Observation: some failures happen because explainers are black-boxes, and both ML practitioners and users do not know their limitations.
The previous argument about interaction failures and the need to adopt a holistic approach of AI systems design also raises the question of the relevance of one-size-fits-all XAI solutions. The most well-known XAI methods (e.g. SHAP and LIME) are often conceived under assumptions of data-, model-, and user-agnosticity (i.e. absence of user focus), generally with the aim of providing more flexibility in their use. Yet, some of the failures discussed in the third section about system-based failures and the previous argument about interaction failures suggest that these methods may not, in fact, be one-size-fits-all solutions. This adds up to the existing research already questioning the often claimed user-agnosticity, suggesting that some types of explanations may in fact not be suited to all user profiles (Wang et al. 2019)), or to all decision models (Alvarez-Melis and Jaakkola 2018b). This restates that explainers face various limitations and various assumptions, that should be known and understood for proper use.
@@
Unfortunately, these limitations and assumptions of XAI systems are rarely known to users and even ML developers, as analyzed in the typology we propose. While this may seem intuitive for user-specific failures (e.g. overtrust issues in biased inferences), it is also true for system-based failures: for instance, some competing issues are caused by the user not knowing how the explainer handles correlations and interactions between features; some instability issues by the user not understanding the tradeoff between stability and locality, nor what level of locality they would like; some incompatibility issues by the user not understanding the differences actually captured by two explainers. Consequently, users may reject or misuse the explanation, not because it does not meet their needs, but because they do not understand how to use it, as shown empirically by  (Kaur et al. 2020) for SHAP.
@@
Technical design choices of XAI systems and the assumptions they rely on heavily impact how the explanation should be interpreted. Overall, this further confirms the need for more effort in communicating on the core capabilities of XAI systems, e.g. through design principles such as ML transparency (Bove et al. 2022). This aligns with prior calls to further encourage more interdisciplinary works for designing more transparent XAI systems, instead of building black-box explainers. However, research in this remains timid and more efforts should be made. One direction to pursue is the possibility of describing explanation methods in a factual way, similarly to Model Cards (Mitchell et al. 2019).
Towards more XUI
Observation: some failures happen because the users are limited in their interaction with the explanations.
From a cognitive point of view, the explanation process is argued to be interactive (Miller 2019) and yet, at best, it is merely sequential in XAI. Most approaches are limited to the generation of factual information about the model’s behaviour or the predicted outcome (e.g., feature importance scores, counterfactual examples, etc.) and the users are not always able to interact with it.

### v2

@@ 5 Proposed Typology: User-specific
5.1 Mismatch
First, we propose to define mismatched explanations when the format of the extracted information does not meet the users’ expectations towards the ML system, leading to the remark ”This is not what I want” in Fig. 2.
@@ 5.1 Mismatch
Depending on the context of the interaction and the nature of the decision model’s outputs, users have been shown to have different needs and questions regarding the ML system, that may also vary depending on the model’s output  (Liao, Gruen, and Miller 2020; van der Waa et al. 2021). In parallel, there is a huge diversity in the forms explanations can take (see e.g. Dwivedi et al. (2023)), but explanation techniques do not necessarily take into account the context in which the explanations is sought by a user (Matarese, Rea, and Sciutti 2021). Therefore, mismatches between the user questions and the generated explanations can occur, on the explanation type (e.g., feature importance or rules), locality (e.g., local or global), goal (e.g., factual or causal) and on the information complexity (e.g., expressed for ML practitioners or lay users), to name a few.
Such mismatches reflect the lack of user-centricity in the conception of the XAI system. Actually, most XAI approaches are designed without evaluating whether the explanations satisfy the needs of real users (Doshi-Velez and Kim 2017; Keane et al. 2021). Instead, other criteria are used to evaluate the relevance of an XAI approach, such as the visual aspect of the explanations, its popularity or the ease of implementation (Mohseni, Zarei, and Ragan 2018; Barredo Arrieta et al. 2020). Such evaluations can bring to light that they fail to meet these needs: for example, during the co-design workshop for an AI-based diagnosis tool (Wang et al. 2019), many interviewed doctors reported that they would prefer alternative hypotheses (e.g., counterfactual examples) rather than factual explanations (e.g., local feature importance scores) that were initially implemented in the system.
@@
Mismatched explanations can lead to dissatisfaction and rejection of the whole ML system: it has been demonstrated that differences between initial expectations and actual experiences can affect both the user satisfaction and acceptance of a system (Bhattacherjee 2001). Along the same lines, the complexity of some explanation types can hinder users to understand the explanations and thus the adoption of the whole ML system (see e.g. Kaur et al. (2020) in the case of SHAP and GAM). It has also been demonstrated that users assign artificial agents human-like traits, and hence expect these agents to provide explanations using the same conceptual framework they are used to (De Graaf and Malle 2017). Yet, XAI solutions have not reached such a complete human-centered approach (Miller 2019; Matarese, Rea, and Sciutti 2021).
@@
Mitigating the mismatched explanations somehow obviously relies on applying user-centered approaches when designing XAI systems: two examples include adjusting the types of explanations according to users’ expectations to improve their satisfaction and acceptance (Riveiro and Thill 2022; Pazzani et al. 2022) or integrating a lexical alignment step to improve the understanding of explanations provided by a conversational agent (Srivastava, Theune, and Catala 2023). Concept-bottleneck models (see e.g. Poeta et al. (2023); Zarlenga et al. (2024)) may be seen as examples of this family of approaches, when checking that the used concepts are meaningful for the users. Another type of solution aims at integrating social and cognitive science’s theories in a theoretical framework for XAI system, e.g. to provide personalized and contextualized explanations (Matarese, Rea, and Sciutti 2021; Byrne 2023).
5.2 Counterintuitive Explanations
We then propose to identify counterintuitive explanations when the explanation provided by the ML system is in contradiction with the prior knowledge of expert users (i.e., AI practitioners and domain experts). Contradiction here occurs at a content level, as opposed to the format level discussed in the case of counterintuitive failures in the previous section.
@@ 5.2 Counterintuitive Explanations
Prior knowledge may take various definitions (Dochy and Alexander 1995), we view it here as ”stored knowledge about the world that have been acquired by an individual” (Brod, Werkle-Bergner, and Shing 2013), including domain expertise and past experience. In an XAI context, users may find faithful explanations contradicting with such prior knowledge, leading them to perceive these explanations as different from what they anticipated (”This is not what I expected” in Fig. 2). This is all the more likely to happen as explanations are especially requested when the model’s output is perceived by users as abnormal or absurd (Riveiro and Thill 2022). In the same vein, past experience has been shown to lead to disagreement with the explanations (Sohn et al. 2019; Suffian et al. 2022).
@@
Counterintuitive explanations do not necessarily represent a failure: explanations that do not match user expectations can indeed be used to fix potential issues within the ML system (Kaur et al. 2020) or for knowledge discovery (Jiménez-Luna, Grisoni, and Schneider 2020). In other situations however, perceiving the provided explanations as counterintuitive can lead users to question the reliability of the prediction even when it is accurate, see e.g. Collaris, Vink, and van Wijk (2018); Palaniyappan Velumani et al. (2022) for studies in applied contexts. The ML system as a whole may be perceived more negatively (Cabitza et al. 2024; Nourani et al. 2021; Ebermann, Selisky, and Weibelzahl 2023), potentially impacting users’ willingness to engage with AI (Ebermann, Selisky, and Weibelzahl 2023). This aligns with findings from works in social sciences, which have shown that people tend to ignore information inconsistent with their beliefs from past experiences (Thagard 1989; Nickerson 1998).
@@
Mitigating these failures requires to better align ML explanations and users’ prior knowledge. Some works in XAI thus argue for more personalized explanations (Ebermann, Selisky, and Weibelzahl 2023; Conati et al. 2021) that would directly integrate in their generation the user knowledge, e.g. expressed as features importance scores (Jeyasothy et al. 2022) or diagrams describing the reasoning process (Lim et al. 2025). Conversely, the predictive model itself can be changed so the generation of explanations that are aligned with users’ prior knowledge is facilitated (Rieger et al. 2020; Koh et al. 2020; Ross, Hughes, and Doshi-Velez 2017). In a different perspective, other works propose to co-design explanation interfaces together with experts so as to integrate both their needs and knowledge (Wang et al. 2019; Weitz et al. 2024).
5.3 Biased Inferences
Finally, we propose to identify explanation failures relying on biased inferences when users make inaccurate interpretations of the explanations, due to cognitive biases. As compared to counterintuitive failures that occur for expert users, for biased inferences we consider mainly lay users who do not have expert knowledge nor past experiences. Yet, such failures can occur for any type of users as cognitive biases are inherent to all humans.
@@ 5.3 Biased Inferences
Similarly to prior beliefs, biases can influence how users respond to different styles of explanations (Liao, Gruen, and Miller 2020; Miller 2019), and several cognitive biases have been shown to trigger inaccurate interpretations of explanations (Bertrand et al. 2022; Nourani et al. 2021; Wang et al. 2019). We discuss below some common cognitive biases with their consequences on the interpretation of explanations.
@@
First, some biases can trigger over reliance in ML explanations. It has been shown that having an explanation, regardless of its quality, increases trust (Hoff and Bashir 2015; Eiband et al. 2019; Lai and Tan 2019). In other examples, it is shown that longer, richer, explanations are found to be more plausible than shorter ones (Fürnkranz, Kliegr, and Paulheim 2020; He, Aishwarya, and Gadiraju 2025), and that users may believe they understand better than what they actually do (Rozenblit and Keil 2002; Mueller et al. 2019; Chromik et al. 2021; He, Kuiper, and Gadiraju 2023). On the other hand, other biases can trigger under reliance. For instance, the ”negativity bias” can cause lay users, in particular, to pay more attention or overweight negative information over positive one of the same strength (Kliegr, Bahník, and Fürnkranz 2021). It may lead users to pay more attention to negative outcomes of the ML system, thus eroding their trust (Pratto and John 1991). It has been demonstrated that showing the weaknesses of the system (e.g., competing explanations) or negative outcomes (e.g., a malignant diagnosis) early on can have a major influence on trust (Nourani et al. 2021). Finally, some other biases can trigger users to misapply the explanations: e.g., the ”insensitivity to sample size” bias may lead lay users to ignore the statistical significance of a statement (Fürnkranz, Kliegr, and Paulheim 2020); the ”availability” may lead lay users to believe that examples and events that easily come to mind are more representative than is actually the case (Wang et al. 2019); e.g., the ”primacy effect” bias may lead them to form an opinion based solely on the first piece of information received (Nourani et al. 2021).
@@
6 Discussion
The typology of explanation failures presented in the previous section allows to understand why failures happen, how to mitigate them, and how to distinguish them from one another. In this section, we leverage this typology to discuss some key issues, and identify promising research directions.
6.1 Towards a Holistic XAI Approach
@@
Many of the explanation failures discussed in the previous sections can be diagnosed as stemming from one of these components (model, explainer and user): they can for instance be due to a technical problem with the system. However, our analysis in the previous section also underlines that some errors actually arise from the interplay between the different components considered, rendering them incompatible: we discussed the interaction between the explainer and the user in Section 5 and discussed in Section 4 some cases of problematic interaction between the model and the explainer. For instance, issues like instability may highlight a mismatch between the decision model’s behavior (volatile, local), the implicit assumptions made during explanation design (stable boundaries), or the similarity between instances perceived by the user.
@@
In the case of these interactions between components, the failure is not caused by a deficiency in either component, but rather from their misalignment. Consequently, some technical solutions suggest adapting or replacing one of the components afterwards, sometimes taking a paradoxical turn: instead of questioning the explainer when observing a failure, it is sometime envisaged instead to train a new AI decision model that would be more adapted to this explainer. As an illustrative example, it has been proposed to build models with more stable behavior to mitigate instability issues of activation-based explainer (Alvarez-Melis and Jaakkola 2018b), or models constrained to minimize the disagreement between LIME and SHAP (Schwarzschild et al. 2023). This may seem surprising, as explainers are usually leveraged to generate insights about the model, not the other way round.
@@
This problem underscores a significant challenge: the interplay between the elements of the system should be taken into account from the system’s inception. The three components of the system should be regarded as interconnected, rather than designed independently. This does not mean abandoning post-hoc methods but rather anticipating their integration in the overall explanation process. Multiple calls for a user-centric approach of XAI have been made, advocating for integrating user needs from the inception of the system and guiding the design of XAI methods (Wang et al. 2019; Ribera and Lapedriza 2019; Vellido 2020; Schmude et al. 2023). Nevertheless, more efforts should be pursued on the interaction between AI models and explainers. One possibility is to draw inspiration from research in Integrative Design for software systems (Tumer and Smidts 2010), proposing holistic design strategies for software systems. Design and monitoring of AI systems should thus be conceived in a holistic way, with any choice or change in the decision model prompting a reassessment of its compatibility with the explainer, and vice versa; and similarly for changes in user requirements.
6.2 Towards More Transparency and Personalization in XAI Systems
Observation: some failures happen because explainers are black-boxes, and both ML practitioners and users ignore their limitations.
The previous argument about the need to adopt a holistic approach of AI systems design also raises the question of the relevance of one-size-fits-all XAI solutions. The most well-known XAI methods (e.g. SHAP and LIME) are often conceived under assumptions of data-, model-, and user-agnosticity (i.e. absence of user focus), generally with the aim of providing more flexibility in their use. Yet, some of the system-specific failures discussed (see Sect. 4) and the previous argument about failures resulting from the interaction of several components suggest that these methods may not, in fact, be one-size-fits-all solutions. This adds up to the existing research questioning this user-agnosticity, suggesting that some types of explanations may not be suited to all user profiles (Hoff and Bashir 2015; Wang et al. 2019), or to all decision models (Alvarez-Melis and Jaakkola 2018b; Molnar et al. 2020) This restates that explainers face various limitations and various assumptions, that should be known and understood for proper use.
@@ 6.2 Towards More Transparency and Personalization in XAI Systems
Unfortunately, these limitations and assumptions of XAI systems are rarely known to users and even ML developers, as analyzed in the proposed typology. While this may seem intuitive for user-specific failures (e.g. overtrust issues in biased inferences), it also holds for system-based failures: for instance, some competing issues are caused by the user not knowing how the explainer handles correlations and interactions between features; some instability issues by their not understanding the tradeoff between stability and locality, nor what level of locality they wish; some incompatibility issues by their not understanding the differences actually captured by two explainers. Consequently, users may reject or misuse the explanation, not because it does not meet their needs, but because they do not understand how to use it, as shown by  Kaur et al. (2020).
@@
Technical design choices of XAI systems and the assumptions they rely on heavily impact how the explanation should be interpreted. Overall, this further confirms the need for more effort in communicating on the core capabilities of XAI systems, e.g. through design principles such as ML transparency (Bove et al. 2022), and more generally improving algorithm literacy of non-ML users (Cabitza, Rasoini, and Gensini 2017; Chiang and Yin 2022; He, Aishwarya, and Gadiraju 2025). This also aligns with prior calls to further encourage interdisciplinary works for designing more transparent XAI systems, instead of building black-box explainers. One possible direction to pursue is improving the standardization of XAI methods (Haque, Islam, and Mikalef 2023), e.g. through the description of explanation methods in a factual way, akin to Model Cards (Mitchell et al. 2019).
6.3 Towards More XUI
Observation: some failures happen because users are limited in their interaction with explanations.
From a cognitive point of view, the explanation process is argued to be interactive (Miller 2019) and yet, at best, it is merely sequential in XAI. Most approaches are limited to the generation of factual information about the model’s behaviour or the predicted outcome (e.g., feature importance scores, counterfactual examples, etc.) and users are often not able to interact with it.